# Slot Abstractors: Toward Scalable Abstract Visual Reasoning

### v1

@@ Abstract
Abstract visual reasoning is a characteristically human ability, allowing the identification of relational patterns that are abstracted away from object features, and the systematic generalization of those patterns to unseen problems. Recent work has demonstrated strong systematic generalization in visual reasoning tasks involving multi-object inputs, through the integration of slot-based methods used for extracting object-centric representations coupled with strong inductive biases for relational abstraction. However, this approach was limited to problems containing a single rule, and was not scalable to visual reasoning problems containing a large number of objects. Other recent work proposed Abstractors, an extension of Transformers that incorporates strong relational inductive biases, thereby inheriting the Transformer’s scalability and multi-head architecture, but it has yet to be demonstrated how this approach might be applied to multi-object visual inputs. Here we combine the strengths of the above approaches and propose Slot Abstractors, an approach to abstract visual reasoning that can be scaled to problems involving a large number of objects and multiple relations among them. The approach displays state-of-the-art performance across four abstract visual reasoning tasks.
@@ 1 Introduction
Specifically, we combine Slot Attention (Locatello et al., 2020), a method for unsupervised object-centric representation learning, with Abstractors (Altabaa et al., 2023). The combined approach, Slot Abstractors (Figure 1), inherits the $O(N^{2})$ complexity of Transformers, allowing it to be scaled to much more complex problems than previous abstract visual reasoning methods. The approach also inherits the Transformer’s multi-head architecture, enabling it to solve problems involving multiple relations. We evaluate the Slot Abstractor on four abstract visual reasoning tasks, with a diverse range of visual and rule complexity. We find that the Slot Abstractor is capable of strong systematic generalization of learned abstract rules, and can be scaled to problems with multiple rule types and more than 100 objects, significantly improving over a number of competitive baselines in most settings.
@@ 2.2 Relational Representation Learning
where $\boldsymbol{W_{q}}^{h}\in\mathbb{R}^{D\times D},\boldsymbol{W_{k}}^{h}\in\mathbb{R}^{D\times D},\boldsymbol{W_{v}}^{h}\in\mathbb{R}^{D\times D}$ are the linear projection matrices used by the $h^{\text{th}}$ head to generate queries, keys, and values respectively; $\boldsymbol{\tilde{s}}^{h}$ is the result of relational cross-attention in the $h^{\text{th}}$ head; and $\boldsymbol{W_{o}}$ are the output weights through which the concatenated outputs of all $H$ heads are passed. The key difference relative to standard cross-attention is that queries and keys are generated from the input features, enabling the relations between those features to be modeled through inner products, and the values are generated from a separate set of (position-specific) embeddings. This relational variant on cross-attention implements a relational bottleneck (Webb et al., 2023a), in that downstream processing (which only has access to $\boldsymbol{s}$) is driven only by the relations between objects, and thus is disentangled from the individual features of those objects. We hypothesized that this relational bottleneck would enable systematic generalization to problems with previously unseen object features.
In each layer, multi-head relational cross-attention was composed together with feedforward networks and standard multi-head self-attention, with residual connections after each of these, as depicted in Figure 1, and described in detail in Algorithm 1. The application of multiple layers of relational and self-attention enabled the Slot Abstractor to flexibly model higher-order relations (relations between relations), which play an important role in abstract visual reasoning. Crucially, this approach retains the quadratic complexity of standard Transformers (given $K$ slots, Slot Abstractors have $O(K^{2})$ complexity), making it feasible to scale the approach to visual inputs containing a large number of objects.
@@ 3 Related Work
A number of neural network models have been proposed (Barrett et al., 2018; Steenbrugge et al., 2018; Zhang et al., 2019b; Zheng et al., 2019; Spratley et al., 2020; Jahrens & Martinetz, 2020; Wang et al., 2020; Wu et al., 2020; Benny et al., 2021; Hu et al., 2021; Zhuo & Kankanhalli, 2022; Zhang et al., 2022) to solve abstract visual reasoning problems (Barrett et al., 2018; Zhang et al., 2019a) based on Raven’s Progressive Matrix (RPM) (Raven, 1938). Many of these models incorporate problem-specific inductive biases, and therefore cannot be applied to other types of visual reasoning problems. To address this, the Slot Transformer Scoring Network (STSN) was proposed (Mondal et al., 2023), combining a transformer reasoning module with a general-purpose inductive bias for object-centric visual processing (Greff et al., 2019; Burgess et al., 2019; Locatello et al., 2020; Engelcke et al., 2021; Dittadi et al., 2021; Jiang et al., 2023). This approach achieved state-of-the-art performance on RPM-based abstract visual reasoning tasks, consistent with results from object-centric models in other visual reasoning tasks (Ding et al., 2021; Wu et al., 2022). However, unlike human abstract reasoning, these approaches generally require very large training sets, and tend to generalize poorly out-of-distribution.
@@ 4.1 Datasets
We evaluated the Slot Abstractor on four challenging abstract visual reasoning datasets, ART (Webb et al., 2021), SVRT (Fleuret et al., 2011), CLEVR-ART (Webb et al., 2023b), and PGM (Barrett et al., 2018). Problems in ART and SVRT datasets consist of simple 2D shapes, whereas CLEVR-ART consists of more complicated and realistic 3D shapes. Each problem of ART, SVRT, and CLEVR-ART contains a small number of objects and is governed by a single rule among them. Problems from the PGM dataset contain a much larger number of objects, with multiple rules among them. The datasets are described in more detail below.
@@ 4.1.1 ART
The Abstract Reasoning Tasks (ART) dataset was proposed by Webb et al (2021). It consists of four visual reasoning tasks (‘same/different’, ‘relational-match-to-sample’, ‘distribution-of-3’, ‘identity rules’), each defined by a different abstract rule (Figure 2). This dataset was created using 100 unicode character objects, and generalization regimes of varying difficulty were created defined by the number of unique objects used to instantiate the rules during training. We focused on the most difficult generalization regime, in which the training set consists of problems that are created using 5 out of the 100 possible objects, and the test problems are created using the remaining 95 objects. This is a difficult test of systematic generalization, as it requires learning of an abstract rule from a very small set of examples (Table 7), with little perceptual overlap between training and test sets. Consistent with previous work (Webb et al., 2023b) we investigated a version of the tasks with multi-object displays and random spatial jitter (random translation of up to 5 pixels in any direction) applied to each object (which has previously been shown to increase task difficulty for some relational architectures (Vaishnav & Serre, 2023)).
@@ 4.2 Baselines
For the PGM dataset, we compared the Slot Abstractor to several state-of-the-art methods reported in Zhang et al. (2022), including their proposed ARII model. Since our main focus was on systematic generalization, for comparison we only included models that were evaluated on at least one OOD regime. We didn’t use any auxiliary information (i.e., rule labels), and hence for a fair comparison we only compared to baselines that were trained without auxiliary loss.
@@ 4.3 Experimental Details
We pre-trained slot attention using a reconstruction (i.e., autoencoding) objective. For ART, SVRT, and CLEVR-ART datasets we used the pre-trained model from Webb et al. (2023b) [^1] https://github.com/Shanka123/OCRA. For PGM, we pre-trained slot attention on the neutral regime training set. Table 8 and Table 9 describe the hyperparameters for the convolutional encoder and the slot-based spatial broadcast decoder respectively. Training hyperparameters are described in Table 10. After pretraining, we selected the model from the epoch with the lowest mean squared error on the validation set. The slot attention parameters were frozen after pretraining.
We resized the images to $128\times 128$ for ART, SVRT, and CLEVR-ART, and resized the image panels (of which there are 9 per problem) to $80\times 80$ for PGM. Pixels were normalized to the range $[-1,1]$. For SVRT, we also applied random horizontal and vertical flips, consistent with prior work (Vaishnav & Serre, 2023; Webb et al., 2023b).
@@
The hyperparameters for the Abstractor module of the Slot Abstractor are described in Table 11. We initialized $\boldsymbol{W_{q}}$ and $\boldsymbol{W_{k}}$ (the linear projection matrices used to generate queries and keys) with the same values, but allowed them to vary during training. To generate the output, we took the mean of the $K$ final relational representations $\boldsymbol{s}^{L}$, and passed this through a task-specific output layer. For the same/different ART task and SVRT tasks, this layer had a single unit and a sigmoid activation, and the model was trained with binary cross entropy loss. For the other ART tasks, and the PGM dataset, each candidate answer choice was inserted into the problem before being fed as input to the model, and the output was a linear layer with a single unit. This output was treated as a score for the answer choice in each input. A softmax activation was then applied over the scores for all answer choices, and the model was trained with cross-entropy loss. For ART and CLEVR-ART tasks, we used a learning rate of $58e-$ 5, batch size of 16, and the number of training epochs are described in Table 12. For SVRT, we used a learning rate of $54e-$ 5, a batch size of 32, and trained for 2000 epochs on each task. We used a single A100 GPU with 80GB memory for training on ART, SVRT, and CLEVR-ART tasks. For training on PGM regimes, we used 4 A100 GPUs with 80GB memory each, a learning rate of $58e-$ 5, and a batch size of 24 per GPU. The number of training epochs for each PGM generalization regime is displayed in Table 13. For optimization, we used the ADAM (Kingma & Ba, 2014) optimizer, and for implementation, we used the PyTorch library (Paszke et al., 2017). The code is available at https://github.com/slotabstractor/slotabstractor.
@@ 5 Results
Table 4 shows the results on the PGM generalization regimes. The Slot Abstractor demonstrated notable improvements on many of the OOD generalization regimes (Intp., H.O.A.P., H.O.T.P., Extp.) with as much as 19% compared to the previous state-of-the-art model (ARII), and an average improvement of around 6.5% over all the regimes. It is also worth noting that OCRA, previously the state-of-the-art model for ART, SVRT, and CLEVR-ART datasets, could not be trained on the PGM dataset due to memory issues. It was not possible to train the model on an A100 GPU with 80GB memory, even when using a batch size of 1 (given $144$ slots, OCRA would need to process $144^{2}\approx 20$k relational embeddings). This highlights the benefits of the Slot Abstractor’s improved complexity ($O(N^{2})$ vs. $O(N^{4})$ in OCRA), allowing it to be effectively applied to problems involving a much larger number of objects.
@@
Table 6: Ablation study on the four tasks of the ART dataset. Results reflect test accuracy averaged over 10 trained networks($\pm$ standard error).
@@
w/o Self Attention 95.26$\pm$0.9 85.41$\pm$2.0 83.01$\pm$1.5 81.12$\pm$1.8
w/o Factorized Reps. 77.24$\pm$2.8 50.02$\pm$0.1 68.97$\pm$1.0 83.26$\pm$1.1
Replace RCA with CA 80.04$\pm$1.4 58.24$\pm$1.5 38.02$\pm$1.2 51.06$\pm$0.4
@@ 5.1 Ablation Study
To understand the importance of each of the Slot Abstractor’s major components, we performed an ablation study using the ART dataset (Table 6). First, we removed Slot Attention, instead dividing the feature map $\boldsymbol{feat}$ into a $4\times 4$ grid (treating each location within the feature map as an ‘object’), and trained the model end to end. We noticed a significant drop in performance for the relational-match-to-sample and same/different tasks. This demonstrates the importance of using object-centric representations. Second, we removed different components from the Abstractor module. In one case, we removed the self-attention and the subsequent feedforward layer. This impaired performance on the relational-match-to-sample, distribution-of-3, and identity rules tasks, where the ablation of self-attention resulted in 4-7% drop in test accuracy. This demonstrates the importance of modeling higher-order relations through self-attention. In another case, we ablated the factorized slot representations, using the standard slot representations instead of separate feature and position embeddings. The relation embeddings $\boldsymbol{s}^{l=0}$ were initialized using learned parameters. We observed a significant drop in performance (as much as 40%) for all tasks except identity rules. This demonstrates the importance of using factorized feature and position embeddings, which allow the Slot Abstractor to keep track of the correspondence between relations and objects. Finally, we replaced the relational cross-attention ($\operatorname{RCA}$) with standard cross-attention ($\operatorname{CA}$), where keys and values are formed from feature embeddings, and queries are formed from position embeddings, thereby removing the relational bottleneck. This significantly impaired performance for all tasks (by as much as 52%). This reflects the centrally important role of the relational bottleneck as an inductive bias underlying the Slot Abstractor’s capacity for generalization.

### v2

@@ Abstract
Abstract visual reasoning is a characteristically human ability, allowing the identification of relational patterns that are abstracted away from object features, and the systematic generalization of those patterns to unseen problems. Recent work has demonstrated strong systematic generalization in visual reasoning tasks involving multi-object inputs, through the integration of slot-based methods used for extracting object-centric representations coupled with strong inductive biases for relational abstraction. However, this approach was limited to problems containing a single rule, and was not scalable to visual reasoning problems containing a large number of objects. Other recent work proposed Abstractors, an extension of Transformers that incorporates strong relational inductive biases, thereby inheriting the Transformer’s scalability and multi-head architecture, but it has yet to be demonstrated how this approach might be applied to multi-object visual inputs. Here we combine the strengths of the above approaches and propose Slot Abstractors, an approach to abstract visual reasoning that can be scaled to problems involving a large number of objects and multiple relations among them. The approach displays state-of-the-art performance across four abstract visual reasoning tasks, as well as an abstract reasoning task involving real-world images.
@@ 1 Introduction
Specifically, we combine Slot Attention (Locatello et al., 2020), a method for unsupervised object-centric representation learning, with Abstractors (Altabaa et al., 2023). The combined approach, Slot Abstractors (Figure 1), inherits the $O(N^{2})$ complexity of Transformers, allowing it to be scaled to much more complex problems than previous abstract visual reasoning methods. The approach also inherits the Transformer’s multi-head architecture, enabling it to solve problems involving multiple relations. We evaluate the Slot Abstractor on five abstract visual reasoning tasks (including a task involving real-world images), with a diverse range of visual and rule complexity. We find that the Slot Abstractor is capable of strong systematic generalization of learned abstract rules, and can be scaled to problems with multiple rule types and more than 100 objects, significantly improving over a number of competitive baselines in most settings.
@@ 2.1 Object-Centric Representation Learning
Although slot attention works well for synthetic images, it is not as effective for real-world images (Seitzer et al., 2022). To extend our approach to problems involving real-world images, we used DINOSAUR (Seitzer et al., 2022), an approach in which slot attention is applied to the features obtained from DINO (Caron et al., 2021) (a large-scale pre-trained vision transformer (ViT) (Dosovitskiy et al., 2020)). This approach has been shown to be an effective method for unsupervised instance segmentation of real-world images. The final attention map is used with the ViT’s positional embeddings to compute position-specific embeddings for each slot as described in Equation 2. The slot embeddings obtained after performing slot attention over the ViT’s features constituted the feature-specific embeddings for each slot.
@@ 2.2 Relational Representation Learning
where $\boldsymbol{W_{q}}^{h}\in\mathbb{R}^{D\times D},\boldsymbol{W_{k}}^{h}\in\mathbb{R}^{D\times D},\boldsymbol{W_{v}}^{h}\in\mathbb{R}^{D\times D}$ are the linear projection matrices used by the $h^{\text{th}}$ head to generate queries, keys, and values respectively; $\boldsymbol{\tilde{s}}^{h}$ is the result of relational cross-attention in the $h^{\text{th}}$ head; and $\boldsymbol{W_{o}}$ are the output weights through which the concatenated outputs of all $H$ heads are passed. The key difference relative to standard cross-attention is that queries and keys are generated from the input features, enabling the relations between those features to be modeled through inner products, and the values are generated from a separate set of (position-specific) embeddings. This relational variant on cross-attention implements a relational bottleneck (Webb et al., 2023a), in that downstream processing (which only has access to $\boldsymbol{s}$) is driven only by the relations between objects, as represented by inner products (see Altabaa and Lafferty (2024) for theoretical considerations regarding the use of inner products to represent relations), and thus is disentangled from the individual features of those objects. We hypothesized that this relational bottleneck would enable systematic generalization to problems with previously unseen object features.
In each layer, multi-head relational cross-attention was composed together with feedforward networks and standard multi-head self-attention, with residual connections after each of these, as depicted in Figure 1, and described in detail in Algorithm 1. The application of multiple layers of relational and self-attention enabled the Slot Abstractor to flexibly model higher-order relations (relations between relations), which play an important role in abstract visual reasoning. Crucially, this approach retains the quadratic complexity of standard Transformers (given $K$ slots, Slot Abstractors have $O(K^{2})$ complexity), making it feasible to scale the approach to visual inputs containing a large number of objects. After $L$ layers, the relational representations were averaged over the $K$ slots, and passed through a task-specific output layer to generate a score $y$. Depending on the task, a score $y$ was generated for each answer choice, and the abstractor module was trained using cross entropy loss. More details can be found in the last paragraph of Section 4.3.
@@ 3 Related Work
A number of neural network models have been proposed (Barrett et al., 2018; Steenbrugge et al., 2018; Zhang et al., 2019b; Zheng et al., 2019; Spratley et al., 2020; Jahrens & Martinetz, 2020; Wang et al., 2020; Wu et al., 2020; Benny et al., 2021; Hu et al., 2021; Zhuo & Kankanhalli, 2022; Zhang et al., 2022; Yang et al., 2023) to solve abstract visual reasoning problems (Barrett et al., 2018; Zhang et al., 2019a; Teney et al., 2020) based on Raven’s Progressive Matrix (RPM) (Raven, 1938). Many of these models incorporate problem-specific inductive biases, and therefore cannot be applied to other types of visual reasoning problems. To address this, the Slot Transformer Scoring Network (STSN) was proposed (Mondal et al., 2023), combining a transformer reasoning module with a general-purpose inductive bias for object-centric visual processing (Greff et al., 2019; Burgess et al., 2019; Locatello et al., 2020; Engelcke et al., 2021; Dittadi et al., 2021; Jiang et al., 2023). This approach achieved state-of-the-art performance on RPM-based abstract visual reasoning tasks, consistent with results from object-centric models in other visual reasoning tasks (Ding et al., 2021; Wu et al., 2022). However, unlike human abstract reasoning, these approaches generally require very large training sets, and tend to generalize poorly out-of-distribution.
@@
Though our proposed model employed previously designed components (slot attention and relational cross-attention), combining these components into an effective architecture necessitated many novel design choices. We have proposed a factorized slot representation, in which the position embeddings are used for the initial relational representations $\boldsymbol{s}^{l=0}$, whereas the feature embeddings are used to compute the keys and queries for the relational cross attention. We also find that relational cross attention was most effective when it was interleaved with self attention and feedforward layers in a particular manner. In Section 5.1, we show through ablation experiments targeted toward the use of the factorized slot representation (Row 4 of Table 6), and the interleaved relational-cross-attention/self-attention (Row 3 of Table 6), that these design choices are essential for obtaining strong performance. Thus, the novel contribution of our work is to identify an effective way for combining object-centric and relational components, yielding an architecture that achieves state-of-the-art performance on challenging abstract reasoning tasks.
@@ 4.1 Datasets
We evaluated the Slot Abstractor on five challenging abstract visual reasoning datasets, ART (Webb et al., 2021), SVRT (Fleuret et al., 2011), CLEVR-ART (Webb et al., 2023b), PGM (Barrett et al., 2018), and V-PROM (Teney et al., 2020). Problems in ART and SVRT datasets consist of simple 2D shapes, whereas CLEVR-ART consists of more complicated and realistic 3D shapes. Each problem of ART, SVRT, and CLEVR-ART contains a small number of objects and is governed by a single rule among them. Problems from the PGM dataset contain a much larger number of objects, with multiple rules among them. V-PROM is a matrix reasoning task (similar to PGM) involving real-world images. The datasets are described in more detail below.
@@ 4.1.1 ART
The Abstract Reasoning Tasks (ART) dataset was proposed by Webb et al. (2021). It consists of four visual reasoning tasks (‘same/different’, ‘relational-match-to-sample’, ‘distribution-of-3’, ‘identity rules’), each defined by a different abstract rule (Figure 3). This dataset was created using 100 unicode character objects, and generalization regimes of varying difficulty were created defined by the number of unique objects used to instantiate the rules during training. We focused on the most difficult generalization regime, in which the training set consists of problems that are created using 5 out of the 100 possible objects, and the test problems are created using the remaining 95 objects. This is a difficult test of systematic generalization, as it requires learning of an abstract rule from a very small set of examples (Table 7), with little perceptual overlap between training and test sets. Consistent with previous work (Webb et al., 2023b) we investigated a version of the tasks with multi-object displays and random spatial jitter (random translation of up to 5 pixels in any direction) applied to each object (which has previously been shown to increase task difficulty for some relational architectures (Vaishnav & Serre, 2023)).
@@ 4.1.4 PGM
4.1.5 V-PROM
The Visual Progressive Matrices (V-PROM) dataset was proposed by Teney et al. (2020). V-PROM is a matrix reasoning dataset (similar to PGM), but unlike standard matrix reasoning tasks, the problems in V-PROM are constructed from real-world images (Figure 7). The maximum number of objects possible in an image is 10, and since there are 9 panels per problem, the total maximum number of objects is 90. Like PGM, each problem consists of multiple abstract rules and the dataset consists of generalization regimes of varying difficulty. In this work, we focused only on the Neutral regime, with around 139K training problems, 8K validation problems, and 73K test problems.
@@ 4.2 Baselines
For the PGM dataset, we compared the Slot Abstractor to the PredRNet model (Yang et al., 2023) and several state-of-the-art methods reported in Zhang et al. (2022), including their proposed ARII model. Since our main focus was on systematic generalization, for comparison we only included models that were evaluated on at least one OOD regime. For the V-PROM dataset, we compared the Slot Abstractor to the best performing model (RN) reported in Teney et al. (2020). We didn’t use any auxiliary information (i.e., rule labels), and hence for a fair comparison we only compared to baselines that were trained without auxiliary loss.
@@ 4.3 Experimental Details
We pre-trained slot attention using a reconstruction (i.e., autoencoding) objective. For ART, SVRT, and CLEVR-ART datasets we used the pre-trained model from Webb et al. (2023b) [^1] https://github.com/Shanka123/OCRA. For PGM and V-PROM, we pre-trained slot attention on the neutral regime training set. Table 8 and Table 9 describe the hyperparameters for the convolutional encoder and the slot-based spatial broadcast decoder respectively, whereas the pre-training hyperparameters are described in Table 12 for the PGM dataset. For the V-PROM dataset, Table 10 describes the hyperparameters for the convolutional encoder used on top of the pre-trained ViT 222we used the features of ViT (token dimensionality 768, 12 heads, patch size 16) pre-trained using DINO, and provided by the timm library (Wightman et al., 2019), with the model name vit_base_patch16_224.dino. features, Table 11 describes the hyperparameters for the decoder, and Table 13 describes the pre-training hyperparameters. After pre-training, we selected the model from the epoch with the lowest mean squared error on the validation set. The slot attention parameters were frozen after pre-training, except for the V-PROM dataset where they were also finetuned on the reasoning task.
We resized the images to $128\times 128$ for ART, SVRT, and CLEVR-ART, and resized the image panels (of which there are 9 per problem) to $80\times 80$ for PGM, and $224\times 224$ for V-PROM. Pixels were normalized to the range $[0,1]$ for ART, and to $[-1,1]$ for SVRT, CLEVR-ART, and PGM datasets. For V-PROM, we applied a channel-wise mean and standard deviation normalization with values of $[0.485,0.456,0.406]$ for the mean, and $[0.229,0.224,0.225]$ for the standard deviation corresponding to the RGB channels, after converting the pixel values to the range $[0,1]$. For SVRT, we also applied random horizontal and vertical flips, consistent with prior work (Vaishnav & Serre, 2023; Webb et al., 2023b).
@@
The hyperparameters for the Abstractor module of the Slot Abstractor are described in Table 14. We initialized $\boldsymbol{W_{q}}$ and $\boldsymbol{W_{k}}$ (the linear projection matrices used to generate queries and keys) with the same values, but allowed them to vary during training. To generate the output, we took the mean of the $K$ final relational representations $\boldsymbol{s}^{L}$, and passed this through a task-specific output layer. For the same/different ART task and SVRT tasks, this layer had a single unit and a sigmoid activation, and the model was trained with binary cross entropy loss. For the other ART tasks, PGM, and the V-PROM datasets, each candidate answer choice was inserted into the problem before being fed as input to the model, and the output was a linear layer with a single unit. This output was treated as a score for the answer choice in each input. A softmax activation was then applied over the scores for all answer choices, and the model was trained with cross-entropy loss. More experimental details can be found at Section A.2. The code is available at https://github.com/Shanka123/Slot-Abstractor.
@@ 5 Results
Table 4 shows the results on the PGM generalization regimes. The Slot Abstractor demonstrated notable improvements on many of the OOD generalization regimes (Intp., H.O.T.P., Extp.) with as much as 21% compared to the previous state-of-the-art model (PredRNet), and an average improvement of around 5% over all the regimes. It is also worth noting that OCRA, previously the state-of-the-art model for ART, SVRT, and CLEVR-ART datasets, could not be trained on the PGM dataset due to memory issues. It was not possible to train the model on an A100 GPU with 80GB memory, even when using a batch size of 1 (given $144$ slots, OCRA would need to process $144^{2}\approx 20$k relational embeddings). This highlights the benefits of the Slot Abstractor’s improved complexity ($O(N^{2})$ vs. $O(N^{4})$ in OCRA), allowing it to be effectively applied to problems involving a much larger number of objects.
On V-PROM, the Slot Abstractor displayed state-of-the-art performance achieving 67.7% test accuracy, compared to 51.2% by the previous best model, Relation Net. This demonstrates that the Slot Abstractor’s systematic generalization capacity can also be extended to abstract visual reasoning problems involving real-world images.
@@
Table 6: Ablation study on tasks of the ART dataset. Results reflect test accuracy averaged over 10 trained networks($\pm$ standard error).
@@
w/o Self Attention 95.26$\pm$0.9 85.41$\pm$2.0 93.83$\pm$0.6 92.58$\pm$0.7
w/o Factorized Reps. 77.24$\pm$2.8 50.02$\pm$0.1 64.4$\pm$1.5 85.08$\pm$1.5
Replace RCA with CA 80.04$\pm$1.4 58.24$\pm$1.5 43.90$\pm$1.3 60.02$\pm$1.0
@@ 5.1 Ablation Study
To understand the importance of each of the Slot Abstractor’s major components, we performed an ablation study using the ART dataset (Table 6). First, we removed Slot Attention, instead dividing the feature map $\boldsymbol{feat}$ into a $4\times 4$ grid (treating each location within the feature map as an ‘object’), and trained the model end to end. We noticed a significant drop in performance for the relational-match-to-sample and same/different tasks. This demonstrates the importance of using object-centric representations (even when pre-trained DINO ViT features are used, as described in A.3). Second, we removed different components from the Abstractor module. In one case, we removed the self-attention and the subsequent feedforward layer. This impaired performance on the relational-match-to-sample, distribution-of-3, and identity rules tasks, where the ablation of self-attention resulted in 4-7% drop in test accuracy. This demonstrates the importance of modeling higher-order relations through self-attention. In another case, we ablated the factorized slot representations, using the standard slot representations instead of separate feature and position embeddings. The relation embeddings $\boldsymbol{s}^{l=0}$ were initialized using learned parameters. We observed a significant drop in performance (as much as 40%) for all tasks except identity rules. This demonstrates the importance of using factorized feature and position embeddings, which allow the Slot Abstractor to keep track of the correspondence between relations and objects. Finally, we replaced the relational cross-attention ($\operatorname{RCA}$) with standard cross-attention ($\operatorname{CA}$), where keys and values are formed from feature embeddings, and queries are formed from position embeddings, thereby removing the relational bottleneck. This significantly impaired performance for all tasks (by as much as 52%). This reflects the centrally important role of the relational bottleneck as an inductive bias underlying the Slot Abstractor’s capacity for generalization.
The removal of slot attention (Row 2 of Table 6) and the replacement of relational cross-attention with standard cross-attention (Row 5 of Table 6), resulted in a significant drop in performance, which demonstrates the importance of combining slot attention with relational cross-attention in Abstractors. The removal of self-attention and the subsequent feedforward layer after relational cross-attention (Row 3 of Table 6) and the use of standard slot representations without factorizing into position and feature specific embeddings (Row 4 of Table 6) explored different design choices of combining slot attention with the relational cross-attention in Abstractors, both of which significantly underperformed compared to the proposed method, thereby demonstrating that these design choices are essential for obtaining strong performance.
5.2 Visualization of Abstractor’s Relation Embeddings
Figure 2: Visualization of Abstractor’s Relation Embeddings. The output of the first head of relational cross attention, after projecting to the first two principal components for 100 examples from the test set of same/different ART task. Two different clusters are formed corresponding to problems with the same and different relation among the objects.
To demonstrate that the Abstractor can capture the different relations present in the tasks, we performed a qualitative analysis of the relational cross-attention head. Specifically, we visualized the output of the relational cross-attention corresponding to the first head ($\boldsymbol{\tilde{s}}^{1}$ in Equation 5) for 100 examples from the test set of same/different ART task, with equal number of examples belonging to same and different category. We applied principal component analysis (PCA) to the output features after taking the mean over the $K$ slots, to reduce it to 2 dimensions. Figure 2 shows that the two-dimensional output form two different clusters corresponding to images belonging to the same and different categories, indicating that the first relational cross-attention head was able to capture the same/different relation.